OpenStudy (anonymous):

Find the height of a right circular cone whose radius is 5 cm and volume is 100π cm3.

OpenStudy (anonymous):

the formula for volume of a cone is: \[\large V=\frac{1}{3}\pi r^2h \] since the radius = 5 and volume = \(\large 100\pi \), plug it in to the formula: \[\large 100\pi =\frac{1}{3}\pi (5)^2h \] now all you gotta do is solve for h, the height.... can you take it from here?

OpenStudy (anonymous):

im lost from here sorry can u show me the rest